# All for One, and One for All: UrbanSyn Dataset, the third Musketeer of Synthetic Driving Scenes

### v1

@@ Abstract
We introduce UrbanSyn, a photorealistic dataset acquired through semi-procedurally generated synthetic urban driving scenarios. Developed using high-quality geometry and materials, UrbanSyn provides pixel-level ground truth, including depth, semantic segmentation, and instance segmentation with object bounding boxes and occlusion degree. It complements GTAV and Synscapes datasets to form what we coin as the ’Three Musketeers’. We demonstrate the value of the Three Musketeers in unsupervised domain adaptation for image semantic segmentation. Results on real-world datasets, Cityscapes, Mapillary Vistas, and BDD100K, establish new benchmarks, largely attributed to UrbanSyn. We make UrbanSyn openly and freely accessible111www.urbansyn.org.
@@ I Introduction
Human data labeling is a major bottleneck for training and validating many deep perception models. To solve vision tasks, datasets of synthetic images have attracted a great deal of interest [3, 4, 1, 2]. They come with many types of automatically generated ground truths (i.e., labels), and, in contrast to human-produced labels, synthetic ones are accurate and consistent across images. However, since deep models trained on synthetic images must later perform in real-world images, the synth-to-real domain gap must be addressed, which remains an open research topic [5, 6, 7]. As argued in other AI sub-fields [8, 9], we must rely on reasonably strong baselines to avoid illusory gains when developing new proposals for a given task. In our case, this means starting with reasonably accurate synthetic-only baseline models from which to bridge the synth-to-real domain gap. The prevailing approach to addressing the synth-to-real gap typically involves focusing on a single synthetic dataset at a time, which leads to less-than-optimal results. It is important to note that when dealing exclusively with real-world labeled images, training complex models on datasets compiled from various sources has been shown to enhance the accuracy as opposed to training them on each source separately. This increase in accuracy is not produced by the larger number of training samples but by the additional diversity derived from the use of different data sources. This has been observed for core visual tasks such as semantic segmentation [10], object detection [11], and depth estimation [12].
@@
Driven by the idea of multi-source approaches, we introduce the UrbanSyn dataset. A collection of highly realistic synthetic images of urban scenarios generated semi-procedurally. UrbanSyn targets computer vision tasks in driving domains, and features labels including pixel-wise instance and semantic masks, depth, and object bounding boxes with occlusion degree. This novel dataset aims at complementing other publicly available synthetic datasets such as GTAV [1] and Synscapes [2], addressing the synth-to-real domain gap from a multi-source perspective.
@@
In this paper, we demonstrate the value of UrbanSyn by addressing semantic segmentation, a challenging task [13] that has proven to be key for the correct performance of many autonomous driving solutions [14, 15]. UrbanSyn brings diversity to urban environments and objects. We show that UrbanSyn gives rise to stronger semantic segmentation baselines than GTAV and Synscapes datasets when used individually. But most importantly, the strongest baselines arise when GTAV, Synscapes, and UrbanSyn are combined. Inspired by Alexandre Dumas, we refer to the combination of these datasets the The Three Musketeers222All synthetic datasets for one, and one for all.. Our experiments make use of modern and relevant approaches, including DeepLab V3+ [16] and SegFormer [17], evaluated in challenging real-world datasets, such as Cityscapes [18], BDD100K [19], and Mapillary Vistas [20]. We also show that our claim remains valid when these approaches are used for synth-to-real semantic self-labeling. To show this, we rely on several synth-to-real unsupervised domain adaptation (UDA) methods. In particular, we use the HRDA [21] and co-training [22] frameworks. We show that adding UrbanSyn boosts semantic segmentation results to the state-of-the-art on synth-to-real UDA for semantic segmentation. On the other hand, we encourage researchers to not only use UrbanSyn for such a task, but also for others requiring labels for instance segmentation, object detection, and depth estimation. Accordingly, we make UrbanSyn publicly available to the research community for free. In this regard, it is also important to remark that, to generate images that can be used for training AI models, the digital assets used to create UrbanSyn have been carefully curated to ensure they are legally usable for that purpose.
@@
The rest of the paper is organized as follows. Section II overviews the most related works, placing ours in context. Section III details the main components involved in the generation of UrbanSyn. Section IV, shows the usefulness of UrbanSyn for synth-to-real UDA in semantic segmentation. Finally, Section V summarizes the work presented in this paper and outlines related future work.
@@ II Related work
A recent trend to increase the realism of the generated images is to employ high-quality 3D assets together with offline Monte Carlo ray-tracing algorithms like path tracing [31, 2]. This was the case of Synscapes [2], in which the gained photorealism came with an extra rendering cost. Another relevant aspect is the authoring cost employed in pursuing realistic layouts. Several works opted to incorporate procedural modeling in their pipelines, e.g., Synscapes incorporated a procedural scene generator, as well as ProcSy [32], which delegated that task on CityEngine [33]. Another alternative consists of automatically synthesizing scene layouts tailored to solve downstream tasks, e.g., semantic segmentation. Meta-Sim [34] aimed at automatically tuning parameters given a target collection of real images in an unsupervised way, while Meta-Sim 2 [35] aims at learning the scene structure in addition to parameters. In both cases, automatically labeled images were generated through a graphics engine sharing similar trade-offs as those presented for most driving simulators.
@@
Using synthetic images to train semantic segmentation models implies addressing the synth-to-real domain gap, which has attracted a significant amount of research during the last years. We advise readers to consult [13] for a comprehensive review. In this context, it is assumed that no human-provided labels are available for real-world images, thus, synth-to-real semantic segmentation becomes a problem of unsupervised domain adaptation (UDA).
In this context, UDA methods are based on two general ideas, taken in isolation or combination. One idea is to perform a synthetic-real domain alignment of the input/features/output of the deep model under training. The other idea is to perform self-labeling (i.e., producing pseudo-labels) of the real-world images for their use in re-training/back-propagation loops. Attending to how coupled to the underlying deep model architecture is the implementation of these ideas, we find model-centric [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 21] and data-centric [49, 50, 51, 52, 53, 54, 22] methods. Model-centric methods have been developed on top of a particular deep architecture (by accessing specific internal layers or modifying its loss function). Data-centric methods treat deep models pretty much as black boxes and mainly focus on assessing and guaranteeing the reliability of the predicted pseudo-labels. Model- and data-centric methods can be complemented with tailored data augmentation techniques [55, 56, 57], with image-to-image translation [58, 59, 60], and with LAB color space alignment which, in line with findings regarding color spaces and robustness of deep models [61], can meaningfully reduce the synth-to-real domain gap [54, 22].
The vast majority of works use Cityscapes as the real-world (target) domain, while SYNTHIA, GTAV, and/or Synscapes are used as synthetic (source) domains. The predominant experimental setting corresponds to using a single synthetic dataset as the source domain. However, some works have shown the convenience of relying on multiple synthetic sources [62, 41, 54, 22] to reduce the synth-to-real domain gap, which is in agreement with findings in the pure real-world setting [10].
@@
We introduce UrbanSyn, a novel dataset for autonomous driving research generated by combining high-quality 3D assets, semi-procedural tools for scene layout generation, and path tracing rendering techniques to produce labeled images with a high degree of photorealism. By design, UrbanSyn is compact and brings differences from GTAV and Synscapes in terms of road layouts, background content, and traffic participants. To show its usefulness, besides providing synthetic-only semantic segmentation baseline results, we use UrbanSyn for running synth-to-real UDA methods. We use the state-of-the-art HRDA [21] to represent model-centric methods, and the state-of-the-art co-training [22] to represent data-centric methods. As the real-world target domain, our experiments not only consider Cityscapes as in most previous literature but also the more challenging BDD100K and Mapillary Vistas datasets. The obtained results confirm that both in isolation and as part of the Three Musketeers dataset, UrbanSyn fosters synth-to-real semantic segmentation.
@@ III UrbanSyn, the new musketeer
This section summarizes different aspects concerning the generation of UrbanSyn, including the construction of our Operational Design Domains (ODDs) and some image rendering details. Overall, as novel real-world datasets are different from previous ones because they were acquired at different real-world areas, using different camera sensors and lighting conditions, such differences also apply to UrbanSyn when compared to GTAV and Synscapes, making them inherently complementary.
@@ III-A Dataset features
The UrbanSyn dataset provides photorealistic color images, paired with their ground-truth depth maps, pixel-level semantic and instance segmentation, as well as 2D object bounding boxes indicating their occlusion degree as a percentage value w.r.t. the whole object within the camera frustum. We adhere to the common classes of Cityscapes as the established de facto standard. Please check Figure 1 to see some image examples and their corresponding labeled information.
UrbanSyn aims to model real-world imperfections affecting both the geometry and the material appearance of its assets. In contrast to GTAV, we utilize unbiased path tracing instead of videogame deferred rendering. When compared to Synscapes, we enrich realism by simulating atmospheric effects like urban air pollution through physics-based light-transport simulation in participating media. In addition, we exploit procedural assets with randomized geometries and materials to enhance the variability of certain assets, i.e.buildings, and we also simulate different types of vehicle lights, all intending to elevate UrbanSyn to a new level of photorealism.
Moreover, UrbanSyn has been meticulously generated to optimize diversity and training value, balancing efficiency and quality and resulting in a more compact dataset than its predecessors. In particular, we showcase distinctive scene configurations, assets, and illumination profiles, while minimizing redundant images as much as possible (e.g., we avoid capturing similar views under similar environment conditions). As a consequence, UrbanSyn consists of images, making it easy to share and store. As we will see in Section IV, despite its compactness UrbanSyn gives rise to the most accurate models for image semantic segmentation.
@@ III-B Data generation pipeline
The designed pipeline to generate UrbanSyn, summarized in Figure 2, allowed us to seamlessly integrate highly realistic and curated scene layouts with procedurally-generated content based on high-quality assets and materials. Similarly to Synscapes or ProcSy, our dataset is mostly procedurally generated. However, our approach uses real maps from GIS sources, i.e., OpenStreetMaps (OSM), to model realistic scenario layouts. The main roads, lanes, and sidewalks are traced over using Mathworks RoadRunner, as well as potential locations of traffic signs or traffic lights. Then, a parameterized in-house builder script parses and instantiates a 3D variation of a given scenario layout inside the Unity game engine. We mainly utilize Unity as our simulation engine to verify plausible spatial occupancy among all assets, e.g., to avoid geometry collisions. The system allows controlling per-class picking distributions in a statistically meaningful manner and supports per-asset weighted probabilities. Our approach assembles these scenarios by parts (layers), sampling from a predefined set of available 3D assets333These assets were developed in-house or acquired with licenses that allow UrbanSyn redistribution under the CC-BY-SA 4.0 license., and automatically instancing procedural variations developed via Unity scripting and SideFX Houdini.
@@ III-C Rendering details
To generate UrbanSyn we decided to employ adaptive sampling, a technique that adaptively dedicates more samples in those regions where the estimated MC variance noise is higher. We established a threshold of samples-per-pixel (spp) to estimate the initial per-pixel variance and let the adaptive sampling algorithm decide where to continue calculating more samples for each pixel. To further reduce noise we relied on OctaneRender’s built-in deep learning-based denoising [66]. This allowed us to significantly reduce the total rendering time per frame, from potentially several hours to just a hundred seconds, all at the expense of trading MC variance noise with some minimal bias introduced by the denoising. The overhead of the denoiser is almost negligible, as it requires only from to seconds per image. Thus, the images of UrbanSyn were rendered and denoised using a maximum of spp. Figure 3 illustrates how renders using spp and denoising could be perceptually very similar to unbiased PT renders with to more spp (e.g.K or K spp).
@@ III-D Operational Design Domains (ODDs)
Using the aforementioned pipeline we can generate different variations and asset distributions within our ODDs. For UrbanSyn we have considered four, inspired by common real traffic layouts (see the top row of Figure 1). To maintain sufficient control over random shuffling, we precomputed a series of intra-layer variations for static elements and conducted random sampling over them. After that, dynamic traffic participants are positioned and later instantiated directly within the rendering engine. This alternative produced enough variations for our purpose of generating single images. The unbiased path tracer then takes the generated scenarios as input and produces the labeled images. Multiple steps are involved before generating the final RGB images, including tone-mapping and gamma transformations to convert them to low-dynamic range (LDR) in the desired RGB color space.

### v2

@@ Abstract
We introduce UrbanSyn, a photorealistic dataset acquired through semi-procedurally generated synthetic urban driving scenarios. Developed using high-quality geometry and materials, UrbanSyn provides pixel-level ground truth, including depth, semantic segmentation, and instance segmentation with object bounding boxes and occlusion degree. It complements GTAV and Synscapes datasets to form what we coin as the ’Three Musketeers’. We demonstrate the value of the Three Musketeers in unsupervised domain adaptation for image semantic segmentation. Results on real-world datasets, Cityscapes, Mapillary Vistas, and BDD100K, establish new benchmarks, largely attributed to UrbanSyn. We make UrbanSyn openly and freely accessible.
@@ I Introduction
Human data labeling is a major bottleneck for training and validating many deep perception models. To solve vision tasks, datasets of synthetic images have attracted a great deal of interest [3, 4, 1, 2]. They come with many types of automatically generated ground truths (i.e., labels), and, in contrast to human-produced labels, synthetic ones are accurate and consistent across images. However, since deep models trained on synthetic images must later perform in real-world images, the synth-to-real domain gap must be addressed, which remains an open research topic [5, 6, 7]. In fact, domain gaps also exist among real-world images captured by different camera sensors. However, synth-to-real domain adaptation is especially relevant because it opens the door to minimizing costly and not fully safe autonomous driving developments in the real world. For instance, in an autonomous driving simulator, we can force desired traffic scenarios that are difficult to observe in the real world by random roaming and train perception models with data captured during such simulations. Research on synth-to-real domain adaptation pursues to develop methods to adapt these models to perform in the real world, preferably in an unsupervised manner. As argued in other AI sub-fields [8, 9], we must rely on reasonably strong baselines to avoid illusory gains when developing new proposals for a given task. In our case, this means starting with reasonably accurate synthetic-only baseline models to bridge the synth-to-real domain gap. The prevailing approach to addressing the synth-to-real gap typically involves focusing on a single synthetic dataset at a time, which leads to less-than-optimal results. It is important to note that when dealing exclusively with real-world labeled images, training complex models on datasets compiled from various sources has been shown to enhance their accuracy compared to training them on each source separately. This increase in accuracy is not produced by the larger number of training samples but by the additional diversity derived from the use of different data sources. This has been observed for core visual tasks such as semantic segmentation [10], object detection [11], and depth estimation [12].
@@
Driven by the idea of multi-source approaches [13, 14, 15, 16], we introduce the UrbanSyn dataset. A collection of highly realistic synthetic images of urban scenarios generated semi-procedurally. UrbanSyn targets computer vision tasks in driving domains and features labels, including pixel-wise instance and semantic masks, depth, and bounding boxes for objects with occlusion degree. This novel dataset aims at complementing other publicly available synthetic datasets such as GTAV [1] and Synscapes [2], addressing the synth-to-real domain gap from a multi-source perspective. To achieve such a complementarity, UrbanSyn is based on unbiased path tracing instead of the video game deferred rendering technique utilized in GTAV. This results in rendering real-world imperfections in a more realistic manner. Moreover, while Synscapes is also more realistic than GTAV, UrbanSyn generation goes beyond by simulating atmospheric effects through physics-based light-transport simulation in participating media. Additionally, UrbanSyn is generated from totally new 3D environments and assets. All this means that UrbanSyn images show scenes different from those in GTAV and Synscapes in terms of content and appearance.
@@
In this paper, we demonstrate the value of UrbanSyn by addressing semantic segmentation, a challenging task [17, 18, 19] that has proven to be key for the correct performance of many autonomous driving solutions [20, 21, 22]. Note that onboard semantic segmentation involves assigning a traffic-relevant semantic class (e.g., Vehicle, Pedestrian, Road, etc.) to each pixel of an image captured for the automatic understanding of the vehicle’s surroundings and so for performing safe driving maneuvers. Deep learning models are the most effective way to accomplish this visual task. Thus, it is essential to gather training images that bring diversity and accurately labeled semantic information. However, obtaining this labeled data is a time-consuming process, taking 60-90 minutes per image due to the manual human labeling involved. UrbanSyn brings accurate pixel-wise semantic labels for onboard images depicting diverse urban environments and objects. When used individually, we show that UrbanSyn gives rise to stronger semantic segmentation baselines than GTAV and Synscapes datasets. But most importantly, the strongest baselines arise when GTAV, Synscapes, and UrbanSyn are combined. Inspired by Alexandre Dumas, we refer to the combination of these datasets as The Three Musketeers 111All synthetic datasets for one, and one for all. (Figure LABEL:fig:short). Our experiments make use of modern and relevant approaches, including DeepLab V3+ [23] and SegFormer [24], evaluated in challenging real-world datasets, such as Cityscapes [25], BDD100K [26], and Mapillary Vistas [27]. We also show that our claim remains valid when these approaches are used for synth-to-real semantic self-labeling. To show this, we rely on several synth-to-real unsupervised domain adaptation (UDA) methods. In particular, we use the HRDA [28] and co-training [16] frameworks. We show that adding UrbanSyn boosts semantic segmentation results to the state-of-the-art on synth-to-real UDA for semantic segmentation. On the other hand, we encourage researchers to not only use UrbanSyn for such a task but also for others requiring labels, for instance, segmentation, object detection, and depth estimation. Accordingly, we make UrbanSyn publicly available to the research community for free. In this regard, it is also important to note that, to generate images that can be used for training AI models, the digital assets used to create UrbanSyn have been carefully curated to ensure they are legally usable for that purpose.
@@
The rest of the paper is organized as follows: Section II overviews the most related works, placing ours in context. Section III details the main components involved in the generation of UrbanSyn. Section IV shows the usefulness of UrbanSyn for synth-to-real UDA in semantic segmentation. Finally, Section V summarizes the work presented in this paper and outlines related future work.
@@ II Related work
A recent trend to increase the realism of the generated images is to employ high-quality 3D assets together with offline Monte Carlo ray-tracing algorithms like path tracing [37, 2]. This was the case of Synscapes [2], in which the gained photorealism came with an extra rendering cost. Another relevant aspect is the authoring cost employed in pursuing realistic layouts. Several works opted to incorporate procedural modeling in their pipelines; e.g., Synscapes incorporated a procedural scene generator, as well as ProcSy [38], which delegated that task to CityEngine [39]. Another alternative consists of automatically synthesizing scene layouts tailored to solve downstream tasks, e.g., semantic segmentation. Meta-Sim [40] aimed at automatically tuning parameters given a target collection of real images in an unsupervised way, while Meta-Sim 2 [41] aims at learning the scene structure in addition to parameters. In both cases, automatically labeled images were generated through a graphics engine sharing trade-offs similar to those presented for most driving simulators.
@@
Using synthetic images to train semantic segmentation models implies addressing the synth-to-real domain gap, which has attracted a significant amount of research in recent years. We advise readers to consult [18] for a comprehensive review. In this context, it is assumed that no human-provided labels are available for real-world images; thus, synth-to-real semantic segmentation becomes a problem of unsupervised domain adaptation (UDA).
In this context, UDA methods are based on two general ideas, taken in isolation or combination. One idea is to perform a synthetic-real domain alignment of the input/features/output of the deep model under training. The other idea is to perform self-labeling (i.e., producing pseudo-labels) of the real-world images for their use in re-training/back-propagation loops. Attending to how coupled to the underlying deep model architecture is the implementation of these ideas, we find model-centric [42, 43, 44, 45, 46, 14, 47, 48, 49, 50, 51, 52, 53, 28, 54] and data-centric [55, 56, 57, 58, 59, 15, 16] methods. Model-centric methods have been developed on top of a particular deep architecture (by accessing specific internal layers or modifying its loss function). Data-centric methods treat deep models pretty much as black boxes and mainly focus on assessing and guaranteeing the reliability of the predicted pseudo-labels. Model- and data-centric methods can be complemented with tailored data augmentation techniques [60, 61, 62], with image-to-image translation [63, 64, 65], and with LAB color space alignment which, in line with findings regarding color spaces and robustness of deep models [66], can meaningfully reduce the synth-to-real domain gap [15, 16].
Most works use Cityscapes as real-world (target) domain, while SYNTHIA, GTAV, and/or Synscapes are used as synthetic (source) domains. The predominant experimental setting corresponds to using a single synthetic dataset as the source domain. However, some works have shown the convenience of relying on multiple synthetic sources [13, 14, 15, 16] to reduce the synth-to-real domain gap, which is in agreement with findings in the pure real-world setting [10].
@@
We introduce UrbanSyn, a novel dataset for autonomous driving research generated by combining high-quality 3D assets, semi-procedural tools for scene layout generation, and path tracing rendering techniques to produce labeled images with a high degree of photorealism. By design, UrbanSyn is compact and brings differences from GTAV and Synscapes regarding road layouts, background content, and traffic participants. To show its usefulness, besides providing synthetic-only semantic segmentation baseline results, we use UrbanSyn for running synth-to-real UDA methods. We use the state-of-the-art HRDA [28] to represent model-centric methods and the state-of-the-art co-training [16] to represent data-centric methods. As the real-world target domain, our experiments not only consider Cityscapes as in most previous literature but also the more challenging BDD100K and Mapillary Vistas datasets. The obtained results confirm that both in isolation and as part of the Three Musketeers dataset, UrbanSyn fosters synth-to-real semantic segmentation.
@@ III UrbanSyn, the new musketeer
This section summarizes different aspects concerning the generation of UrbanSyn, including the construction of our Operational Design Domains (ODDs) and some image rendering details. Overall, as novel real-world datasets are different from previous ones because they were acquired in different real-world areas using different camera sensors and lighting conditions, such differences also apply to UrbanSyn compared to GTAV and Synscapes, making them inherently complementary.
@@ III-A Dataset features
The UrbanSyn dataset provides photorealistic color images, paired with their ground-truth depth maps, pixel-level semantics and instance segmentation, as well as 2D object bounding boxes indicating their occlusion degree as a percentage value w.r.t. the whole object within the camera frustum. We adhere to the common classes of Cityscapes as the established de facto standard. Please check Figure 1 to see some image examples and their corresponding labeled information.
UrbanSyn aims to model real-world imperfections affecting both the geometry and the material appearance of its assets. In contrast to GTAV, we utilize unbiased path tracing instead of videogame deferred rendering. When compared to Synscapes, we enrich realism by simulating atmospheric effects like urban air pollution through physics-based light-transport simulation in participating media. In addition, we exploit procedural assets with randomized geometries and materials to enhance the variability of certain assets (e.g., buildings), and we also simulate different types of vehicle lights, all intending to elevate UrbanSyn to a new level of photorealism.
Moreover, UrbanSyn has been meticulously generated to optimize diversity and training value, balancing efficiency and quality and resulting in a more compact dataset than its predecessors. In particular, we showcase distinctive scene configurations, assets, and illumination profiles while minimizing redundant images as much as possible (e.g., we avoid capturing similar views under similar environment conditions). As a consequence, UrbanSyn consists of images, making it easy to share and store and being aligned with the train on less data tip for decarbonizing AI development [67]. As we will see in Section IV, despite its compactness, UrbanSyn gives rise to the most accurate models for image semantic segmentation.
@@ III-B Data generation pipeline
The designed pipeline to generate UrbanSyn, summarized in Figure 2, allowed us to seamlessly integrate highly realistic and curated scene layouts with procedurally-generated content based on high-quality assets and materials. Similarly to Synscapes or ProcSy, our dataset is mostly procedurally generated. However, our approach uses real maps from GIS sources, i.e., OpenStreetMaps (OSM), to model realistic scenario layouts. The main roads, lanes, and sidewalks are traced over using Mathworks RoadRunner, as well as potential locations of traffic signs or traffic lights. Then, a parameterized in-house builder script parses and instantiates a 3D variation of a given scenario layout inside the Unity game engine. We mainly utilize Unity as our simulation engine to verify plausible spatial occupancy among all assets, e.g., to avoid geometry collisions. The system allows controlling per-class picking distributions in a statistically meaningful manner and supports per-asset weighted probabilities. Our approach assembles these scenarios by parts (layers), sampling from a predefined set of available 3D assets222These assets were developed in-house or acquired with licenses that allow UrbanSyn redistribution under the CC-BY-SA 4.0 license., and automatically instancing procedural variations developed via Unity scripting and SideFX Houdini.
@@ III-C Rendering details
To generate UrbanSyn, we decided to employ adaptive sampling, a technique that adaptively dedicates more samples in those regions where the estimated MC variance noise is higher. We established a threshold of samples-per-pixel (spp) to estimate the initial per-pixel variance and let the adaptive sampling algorithm decide where to continue calculating more samples for each pixel. To further reduce noise, we relied on OctaneRender’s built-in deep learning-based denoising [71]. This allowed us to significantly reduce the total rendering time per frame, from potentially several hours to just a hundred seconds, all at the expense of trading MC variance noise with some minimal bias introduced by the denoising. The overhead of the denoiser is almost negligible, as it requires only from to seconds per image. Thus, the images of UrbanSyn were rendered and denoised using a maximum of spp. Figure 3 illustrates how renders using spp and denoising could be perceptually very similar to unbiased PT renders with to more spp (e.g.K or K spp).
@@ III-D Operational Design Domains (ODDs)
Using the aforementioned pipeline, we can generate different variations and asset distributions within our ODDs. For UrbanSyn, we have considered four, inspired by common real traffic layouts (see the top row of Figure 1). Table I details the number of images, average, and total rendering time for each ODD. We can observe that using a participating medium has a huge impact on render times. Additionally, Table II details the intrinsic variability of the ODDs, detailing the number of assets, materials, and variations for each class. We build the scenes used for rendering UrbanSyn, combining these assets to obtain the greatest amount of diversity while preserving the plausibility of the result. To maintain sufficient control over random shuffling, we precomputed a series of intra-layer variations for static elements and conducted random sampling over them. After that, dynamic traffic participants are positioned and later instantiated directly within the rendering engine. This alternative produced enough variations for our purpose of generating single images.
The unbiased path tracer then takes the generated scenarios as input and produces the labeled images. Multiple steps are involved before generating the final RGB images, including tone-mapping and gamma transformations, to convert them to low-dynamic range (LDR) in the desired RGB color space.
#ODD |
ODD 1 | 1,319 | No | 36.96s | 13h 32m | |||||||||
ODD 2 | 1,369 | Yes | 120.82s | 45h 56m | |||||||||
ODD 3 | 2,905 | No | 24.55s | 19h 48m | |||||||||
ODD 4 | 1,946 | Yes | 87.95s | 47h 32m | |||||||||
Total | 7,539 | Y:43%, N:57% | 69,17s | 132h 46m |